Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name E.Chudakov
Log entry time 23:57:38 on November 24, 2008
Entry number 250615
keyword=Experiment summary for Nov 23(swing) - 24(day)
Summary for Nov 23 (swing) - Nov 24 (day)
Smooth data taking
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/11/23 swing | 4.89 |
0.28 | 2.83 |
2008/11/24 owl | 6.17 |
0.75 | 1.08 |
2008/11/24 day | 6.20 |
0.27 | 1.53 |
total | 17.26 |
1.30 | 5.44 |
Main events/issues:
- Monday swing: hacsbc2 X-session became hardly usable. I seems
there is an erratic signal associated with the mouse. Neither
replacing
the mouse nor restarting X helped.
Plans for the next few days:
- Running with minimal changes till Tuesday morning;
- Tuesday 8:30 - going to restricted access till about 15:00.
After that we may go to control access in order
to do laser alignment.
The main work scheduled:
- Target cell replacement, survey after replacement,
- Compressor for the dehumidifier - by accelerator people
- Compton ionis. chamber - by accelerator people
- Other work by us: Compton electronic noise, RICH, gas
Cher. etc